# SECap: Speech Emotion Captioning with Large Language Model

### 19 Dec 2023 v2

@@ Introduction
where $W_{q_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{k}}$, $W_{k_{\text{self}}}\in\mathbb{R}^{d_{s}\times d_{k}}$, and $W_{v_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{v}}$ are the learnable weight matrices for queries, keys, and values in the self-attention mechanism, and $d_{k}$ and $d_{v}$ are the dimensions of keys and values. The output of the self-attention mechanism $A_{\text{self}}\in\mathbb{R}^{n_{q}\times d_{v}}$ are then used as the queries for the cross-attention mechanism, while the speech embedding $S\in\mathbb{R}^{n_{s}\times T_{s}\times d_{s}}$ serve as keys and values:
@@
This approach enables the attention mechanism to retrieve features related to Q-queries within the speech embedding. Specifically, the output of the Q-Former, denoted as the Q-Embedding $Q_{e}\in\mathbb{R}^{n_{s}\times d_{q}}$, maintains a fixed length that is independent of the length of the input speech. This fixed-length representation leads to improved generalization performance across speech inputs of varying lengths.
@@
$\mathcal{L}(Q_{c};Q_{e})=\sum_{i=1}^{NK}\bigg{[}w_{1}\mathcal{S}(e_{i},d_{i})+w_{2}\sum_{j=1}^{K-1}\mathcal{S}(e_{i},p_{ij})$
$+w_{3}\sum_{j=1}^{NK-K}\text{ReLU}(\mathcal{S}(e_{i},u_{ij})-m)\bigg{]}$ (5)

### 23 Dec 2023 v3

@@ Introduction
where $W_{q_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{k}}$, $W_{k_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{k}}$, and $W_{v_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{v}}$ are the learnable weight matrices for queries, keys, and values in the self-attention mechanism, and $d_{k}$ and $d_{v}$ are the dimensions of keys and values. The output of the self-attention mechanism $A_{\text{self}}\in\mathbb{R}^{n_{q}\times d_{v}}$ are then used as the queries for the cross-attention mechanism, while the speech embedding $S\in\mathbb{R}^{n_{s}\times T_{s}\times d_{s}}$ serve as keys and values:
@@
This approach enables the attention mechanism to retrieve features related to Q-queries within the speech embedding. Specifically, the output of the Q-Former, denoted as the Q-Embedding $Q_{e}\in\mathbb{R}^{n_{s}\times n_{q}\times d_{q}}$, maintains a fixed length that is independent of the length of the input speech. This fixed-length representation leads to improved generalization performance across speech inputs of varying lengths.
@@
$\mathcal{L}(Q_{c};Q_{e})=\sum_{i=1}^{NK}\Big{[}w_{1}(1-\mathcal{S}(e_{i},d_{i}))+w_{2}\sum_{j=1}^{K-1}(1-$
$&\mathcal{S}(e_{i},p_{ij}))+w_{3}\sum_{j=1}^{NK-K}\text{ReLU}(\mathcal{S}(e_{i},u_{ij})-m)\Big{]}\end{aligned}$ (5)